OpenStudy (anonymous):

A juice company decides to test five different brands of juice. The different brands have been labeled A, B, C, D, and E. The company decides to compare each brand with the other brands by pairing together different brands. How many different pairs will result by selecting two different brands at the same time?

OpenStudy (anonymous):

@OregonDuck

OpenStudy (anonymous):

11 15 120 10

OpenStudy (anonymous):

I am thinking 10 but IDK

oregonduck (oregonduck):

you are right

oregonduck (oregonduck):

you would want to find the number of combinations of 5 things taken 2 at a time 5C2 = 5!/[2!3!] = 5 * 4 * 3 * 2 * 1/(2 * 1 * 3 * 2 * 1) = (5 * 4)/2 = 5 * 2 = 10

OpenStudy (anonymous):

Yay! OK can you do one more?

oregonduck (oregonduck):

yup

OpenStudy (anonymous):

There are 9 candidates running for 3 seats on a committee. How many different election results are possible?

OpenStudy (anonymous):

36 84 504 56

OpenStudy (anonymous):

I think 84

oregonduck (oregonduck):

yup you are right 84 it means 9C3 = 9! / (9-3)!3! = 9! / 6!3! = 9*8*7*6! / 6!*3*2*1 = 84

OpenStudy (anonymous):

Yay! Thank you! You've been a big help!

oregonduck (oregonduck):

fyi you should fan me if you need any more help that way you can send me a message

oregonduck (oregonduck):

yup np anytime

OpenStudy (anonymous):

lol yeah I already did

oregonduck (oregonduck):

oh ok good :)

OpenStudy (anonymous):

Thanks for the help! :)

oregonduck (oregonduck):

yup, anytime you need help just send me a message and i will try to go on openstudy